Session: Feb 2017 (session 1)
Wavelength: 1.3 cm


To give feedback on this experiment, click here

Experiment Feedback

Yebes (Miguel Sanz at Wed Mar 08 04:15:12 2017) Success:

Medicina (Beppe at Wed Mar 08 04:47:02 2017) Success:

Yebes (Daniel Valero at Wed Mar 08 07:46:56 2017) Minor failures: RECO
Half of the channels for Q band have not been recorded (ifb and ifd). After switching back to K band dbbc firmware didn't work properly at bbc1 and bbc3

Robledo (Cristina Garcia Miro at Wed Mar 08 08:14:39 2017) Failure: ANT
After "Follow the Sun" DSN installation and testing antenna controller was declared red. Not possible to observe.

Hartebeesthoek (Jonathan Quick at Wed Mar 08 08:17:20 2017) Success:
No known problems.

Effelsberg (Uwe at Wed Mar 08 08:50:33 2017) Minor failures: FORM
Faulty 10 MHz distributer caused sync problems at the DBBC. Obviously affected time ranges are 13:10 to 14 UT and 16 to 17 UT (wrong gps-fmout measurements), but it is not clear how much the VLBI data suffered from that. No observations between 17 and 19:30 UT to investigate the problem and solve it. Observations continued and went fine after 19:30 UT.

Urumqi (Peng Li at Wed Mar 08 09:56:05 2017) Minor failures:
Because the goddess of the moon observation, UT01:40 stop observation.

Zelenchukskaya (Ptitcin A. at Wed Mar 08 12:07:18 2017) Success:
Scans no0097 - no0140 missed due to participation in intensive geo session

Badary (Lysakova, Nekrasov at Wed Mar 08 12:08:43 2017) Success:
Missed scans No0101 to No0139 - pause for observing intensive geo session

Svetloe (Shumilova at Wed Mar 08 12:09:04 2017) Success:

Onsala (Jun Yang at Wed Mar 08 14:04:21 2017) Success:

Effelsberg (Uwe at Fri Mar 10 07:56:05 2017) Success:
Some additional notes on the weather: rainy and windy through most of the observations and therefore increased Tsys.

Noto (P.Cassaro at Sat Mar 11 10:47:08 2017) Success:

Torun (Pawel Wolak at Fri Mar 17 11:10:38 2017) Success: